OpenStudy (anonymous):

Andy has 28 coins made up of quarters and half dollars, and their total value is $11.50. How many quarters does he have ?

OpenStudy (aaronq):

So you've seen how i've written two equations to solve for these variables. Can you try it out?

OpenStudy (anonymous):

ok

OpenStudy (aaronq):

dont take it in a bad way, if you dont learn for yourself then i'm of no help to you.

OpenStudy (anonymous):

i know . 28 coins in all: x+y=28 amounting to $11.50 :

OpenStudy (anonymous):

0.25x+0.50y=11.50

OpenStudy (aaronq):

yep, thats it! now solve them

OpenStudy (aaronq):

solve for x

OpenStudy (anonymous):

would it be 20 ?

OpenStudy (aaronq):

nope

OpenStudy (aaronq):

0.25x+0.50(28-x)=11.50

OpenStudy (anonymous):

that answer isnt there only 10

OpenStudy (anonymous):

and 20

OpenStudy (aaronq):

i dont understand what you're trying to say. the number of quarters (when you solve the equation for x) is 10.

OpenStudy (anonymous):

okay i get it now sorry

OpenStudy (aaronq):

no worries, what you said just confused me lol